Hi just need some help on these word problems,
trent can deliver his newspaper in 10 minutes. It takes Lois 40 minutes to do the same route. How long would it take if they worked together?
thnks Kim

Found 2 solutions by mananth, richwmiller:
Answer by mananth(16949)

trent takes 10 minutes to distribute papers. ( this we consider as 1 job)
so in 60 minutes he can do 6jobs
.
Lois does same job in 40 minutes.
In 60 minutes he does 60/40 of the job
=3/2 of the job.
.
Both together will do 6 + 3/2 jobs in 1 hour
= 15/2 jobs.
15/2 jobs they do in 1 hour.
1 job they do in how many hours?
.
2/15 hour they take to do one job.
2/15 *60 = 8 minutes

Answer by richwmiller(17219)

x/10+x/40=1
4x+x=40
5x=40
x=8 minutes